OpenStudy (anonymous):

The ratio of the perimeters of two similar triangles is 4:3. What are the areas of these triangles if the sum of their areas is 130cm^2?

OpenStudy (anonymous):

@Zarkon

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

for similar triangles the ratio of the areas is equal to ratio of the squares of their sides

OpenStudy (anonymous):

here the ratio of the sides is 4:3 ( ratio of corresponding sides = ratio of perimeters) thus their areas is in the ratio of areas is 16:9

OpenStudy (anonymous):

thus area of one of the triangle =16/(16+9) * 130 =83.2 and that for the other triangle is 9/(16+9) *130 =46.8

OpenStudy (whpalmer4):

If you're wondering why the ratio of the areas is the ratio of the squares of the sides, consider that the area is \(\dfrac{1}{2}bh\), and so we will be multiplying both \(b\) and \(h\) by the scale factor.